OpenStudy (anonymous):

solve x^2+7x-10=0, find the intercepts of x

OpenStudy (anonymous):

doesn't factorise, so use quadratic formula

OpenStudy (amistre64):

(x^2 +7x +(49/4))-10-(49/4) = 0 (x+(7/2))^2 -89/4 = 0 x+(7/2) = +- sqrt(89)/2 x = -(7/2) +- sqrt(89)/2 maybe

OpenStudy (amistre64):

\[x=\frac{-1\pm \sqrt{89}}{2}\]

OpenStudy (anonymous):

if your job is to complete the square start \[x^2+7x=10\] \[(x+\frac{7}{2})^2=10+(\frac{7}{2})^2\] \[

OpenStudy (anonymous):

then \[(x+\frac{7}{2})^2=10+\frac{49}{4}=\frac{89}{4}\]

OpenStudy (anonymous):

then \[x+\frac{7}{2}=\pm\frac{\sqrt{89}}{2}\]

OpenStudy (anonymous):

and finally subtract to get \[x=-\frac{7}{2}\pm\frac{\sqrt{89}}{2}\] aka \[x=\frac{-7\pm\sqrt{89}}{2}\]

OpenStudy (anonymous):

just like amistre got